# Visual Semantic Navigation with Real Robots

### v1

@@ 4.4 Discussion
Analyzing the particular behavior of the models, we can provide the following interesting discussion. The integration of the VLV and PIRLNav models within our ROS4VSN architecture has proven to be successful. It has resulted in a mobile agent capable of navigating in closed environments autonomously, obtaining numerous experiments where the robotic platforms reach the target class without complications and following logical and direct trajectories. This navigation is comparable to that observed within simulated environments. For the VLV model that integrates an object detector, we have observed that this fact has a significant impact on the agent’s navigation, especially when it is close to the target class. Although the object detector does not significantly affect the general exploration, its impact becomes crucial when the robot is in the vicinity of the target. At this crucial stage of navigation, the object detector provides a significant advantage by guiding the robot to the target more effectively. This explains the difference of performance we have observed between VLV and PIRLNav in the real world.
@@
In terms of qualitative aspects of navigation, we believe that the PIRLNav model is better than the VLV model. Note that VLV, every time the high-level policy has to make a decision, needs the robot to turn completely on itself, taking 12 captures on which it will decide which direction to move forward. This can be observed in the provided video. This feature slows down navigation, although it could be solved with some specific hardware. In contrast, PIRLNav offers a more direct navigation experience. It is interesting to observe the type of action sampling that PIRLNav performs while being executed on our robots. Figure 12 shows a histogram corresponding to the distribution of navigation actions performed by PIRLNav. First, one can observe that the look_up and look_down actions have hardly been selected. This allows us to affirm that the impact of the adaptations we have made to replace these actions by backward and forward movements, respectively, could hardly have had a considerable impact on the final results. Second, the most popular actions, as they are the ones that motivate the exploration of the environment, are those of advances and turns. The stop action was sampled 31 times, that is, a $0.2\%$ as it reflects the histogram provided. We believe that work should be done on solutions to increase the number of times the stop action is selected, but to do so reliably.
Finally, our study confirms some of the conclusions reported in recent works, e.g. (Gervet et al., 2022). Modular-learning models, such as VLV, perform better than end-to-end learning approaches, e.g. PIRLNav, when tested in the real world.
@@ 5 Conclusions
To conclude, we have presented a ROS-based framework for visual semantic navigation named ROS4VSN that allows to easily test and compare different VSN models in real robots. Using ROS4VSN, we have been able to embed two cutting-edge VSN models into two distinct real robotic platforms. The chosen models are PIRLNav (Ramrakhya et al., 2023) and VLV (Chang et al., 2020). To seamlessly integrate these models, technical modifications have been needed. These adaptations ensure a smooth transition for the models, enabling them to shift from interacting with observations generated in simulation environments to those obtained from the real world. We have also offered a thorough experimental evaluation to showcase how these VSN approaches behave when navigating in the real world. Our novel framework shows a robust stability, being able to run for a considerable distance, in two different robots, without any human intervention. Our study and results show that the performance of state-of-the-art VSN models is significantly lower in the real world than in the virtual environments where they were trained. We anticipate that our efforts will lay the foundation for addressing this substantial challenge.
@@ Acknowledgements
This research was funded by project AIRPLANE, with reference PID2019-104323RB-C3 from the Ministry of Science and Innovation of Spain.

### v2

@@ 4.4 Discussion
Analyzing the particular behavior of the models, we can provide the following interesting discussion. The integration of the VLV and PIRLNav models within our ROS4VSN architecture has proven to be successful. It has resulted in a mobile agent capable of navigating in closed environments autonomously, obtaining many experiments where the robotic platforms reach the target class without complications and following logical and direct trajectories. This navigation is comparable to that observed within simulated environments. For the VLV model that integrates an object detector, we have observed that this fact has a significant impact on the agent’s navigation, especially when it is close to the target class. Although the object detector does not significantly affect the general exploration, its impact becomes crucial when the robot is in the vicinity of the target. At this crucial stage of navigation, the object detector provides a significant advantage by guiding the robot to the target more effectively. This explains the difference of performance we have observed between VLV and PIRLNav in the real world.
@@
In terms of qualitative aspects of navigation, we believe that the PIRLNav model is better than the VLV model. Note that VLV, every time the high-level policy has to make a decision, needs the robot to turn completely on itself, taking 12 captures on which it will decide which direction to move forward. This can be observed in the provided video. This feature slows down navigation, although it could be solved with some specific hardware. In contrast, PIRLNav offers a more direct navigation experience. It is interesting to observe the type of action sampling that PIRLNav performs while being executed on our robots. Figure 12 shows a histogram corresponding to the distribution of navigation actions performed by PIRLNav. First, one can observe that the look_up and look_down actions have hardly been selected. This allows us to affirm that the impact of the adaptations we have made to replace these actions by backward and forward movements, respectively, could hardly have had a considerable impact on the final results. Second, the most popular actions, as they are the ones that motivate the exploration of the environment, are those of advances and turns. The stop action was sampled 31 times. This is a $0.2\%$ as it is reflected in the histogram provided. We believe that work should be done on solutions to increase the number of times the stop action is selected, but to do so reliably.
Finally, our study confirms some of the conclusions reported in recent works, e.g [51]. Modular-learning models, such as VLV, perform better than end-to-end learning approaches, e.g PIRLNav, when tested in the real world.
@@ 5 Conclusions
To conclude, we have presented a ROS-based framework for visual semantic navigation named ROS4VSN that allows to easily test and compare different VSN models in real robots. Using ROS4VSN, we have been able to embed two cutting-edge VSN models into two distinct real robotic platforms. The chosen models are PIRLNav [1] and VLV [3]. To seamlessly integrate these models, technical modifications have been needed. These adaptations ensure a smooth transition for the models, enabling them to shift from interacting with observations generated in simulation environments to those obtained from the real world. We have also offered a thorough experimental evaluation to showcase how these VSN approaches behave when navigating in the real world. Our novel framework shows a robust stability, being able to run for a considerable distance, in two different robots, without any human intervention. Our study and results show that the performance of state-of-the-art VSN models is significantly lower in the real world than in the virtual environments where they were trained. We expect that our efforts will lay the foundation for addressing this significant challenge.
@@ Acknowledgements
This research was partially funded by projects: NAVISOCIAL, with reference 2023/00405/001 from the University of Alcalá; NAVIGATOR-D, with reference PID2023-148310OB-I00 from the Ministry of Science and Innovation of Spain.